# A Cross-Font Image Retrieval Network for Recognizing Undeciphered Oracle Bone Inscriptions

### v1

@@ Abstract
Oracle Bone Inscription (OBI) is the earliest mature writing system known in China to date, which represents a crucial stage in the development of hieroglyphs. Nevertheless, the substantial quantity of undeciphered OBI characters continues to pose a persistent challenge for scholars, while conventional methods of ancient script research are both time-consuming and labor-intensive. In this paper, we propose a cross-font image retrieval network (CFIRN) to decipher OBI characters by establishing associations between OBI characters and other script forms, simulating the interpretive behavior of paleography scholars. Concretely, our network employs a siamese framework to extract deep features from character images of various fonts, fully exploring structure clues with different resolution by designed multiscale feature integration (MFI) module and multiscale refinement classifier (MRC). Extensive experiments on three challenging cross-font image retrieval datasets demonstrate that, given undeciphered OBI characters, our CFIRN can effectively achieve accurate matches with characters from other gallery fonts.
@@ I Introduction
Oracle Bone Inscription (OBI) derives its name from being carved on tortoise shells and animal bones, serving as significant records for divination and military warfare during China’s Shang Dynasty. Therefore, deciphering OBI characters enables scholars to access valuable information recorded thousands of years ago, which is crucial for understanding the social structure and cultural practices of early China. To decode undeciphered OBI characters, paleography scholars not only need extensive linguistic knowledge and practical experience, but also must study its historical formations and diverse variations by referring to a large number of ancient books [1]. Due to the intensive time and manpower required in the interpreting process, only about 2,200 of the nearly 4,500 OBI characters discovered have been deciphered so far [2].
Figure 1: The workflow illustration of the proposed CFIRN. We retrieve the most similar character in gallery font annotated with modern Chinese character for an input OBI character.
Figure 2: Architecture of the proposed CFIRN. We extract the high-quality feature representation of input image group $\{\mathbf{OI^{1}},\mathbf{BI^{1}}\}$ trough the OBI Branch and the Gallery Font Branch, and determine if OBI character and gallery font character belong to the same category by comparing the similarity of final feature. In the testing phase, we retrieval the most similar gallery font character to obtain the decryption for each undeciphered OBI character.
@@
With the significant advancements in artificial intelligence technology in recent years, a growing number of researchers have been applying advanced AI methods to tackle various practical challenges in the OBI domain, such as oracle bone character fragments rejoining [3, 4], OBI character detection [5, 6, 7, 8, 9], and OBI character recognition [10, 11, 12, 13, 14, 15, 16, 17, 18], to name a few.
@@
In the field of OBI character decipherment, significant strides have been made through various deep learning-driven approaches. Wang et al.[19] analyzed the evolution process using few-shot learning, while others have explored classification approaches to associate undeciphered OBI with ancient scripts such as the large seal script [20]. The use of diffusion models has also been emerged as a promising avenue for evolving OBI characters into modern forms [21, 22]. Furthermore, innovative methods involving the disassembly and reconstruction of OBI characters into modern substitutes have been proposed [23]. Collectively, these studies underscore the potential of deep learning in advancing the automated decipherment and evolution of ancient scripts.
@@
However, the significant evolution of character forms throughout an extensive historical period presents an impediment to establishing direct and immediate correlations. In the concrete practice of paleography, researchers often employ comparative analysis of OBI characters with scripts from various historical periods, particularly those from closely related epochs, to discern more plausible interpretations.
@@
Drawing insights from the methodology of paleographic research, we design the retrieval-based workflow to decrypt the undeciphered OBI characters as shown in Fig. 1. Concretely, we propose a cross-font image retrieval network (CFIRN) “that establishes associations between OBI characters and other script forms, with the process depicted in Fig. 2. We seamlessly combine ConvNeXt-based encoder features through the application of multiscale feature integration (MFI) module, further enhancing them with multiscale refinement classifier (MRC) to achieve the generation of highly refined features. Consequently, three richly annotated ancient scripts, including Bronze Inscription (BI), Bamboo Slip Inscription (BSI) and Clerical Script (CS), are selected as our gallery set.
@@
In summary, our main contributions can be summarized as follows: 1) To the best of our knowledge, this is the first study to explore the capability of a unified cross-font image retrieval network for the task of OBI character decipherment. 2) Our CFIRN framework introduces a novel multiscale feature integration (MFI) module and multiscale refinement classifier (MRC), enhancing the extraction of scale-specific and subtle semantic information. 3) Compared to state-of-the-art image retrieval models, our model demonstrates effectiveness on the three ancient script image retrieval datasets.
@@ II METHOD
The overall architecture of the proposed CFIRN is illustrated in Fig. 2, which is designed to integrate the multiscale deep information by a siamese network from character images of two fronts, culminating in an enhanced high-quality feature representation. The OBI branch and the Gallery Font branch share a common structural framework, hence we illustrate the OBI branch as a representative example.
@@
II-A Feature Extraction Process
Drawing inspiration from related research [24, 25], we adopted ConvNeXt [26] as the backbone of the encoder network. In detail, we implement the encoder weight sharing between the OBI branch and the Gallery Font branch, enhancing the generalization of the encoder in extracting character image features under low-data conditions. Given an input image set $\{\mathbf{OI^{1}},\mathbf{BI^{1}}\}$, including an OBI character image and another gallery font character image, we obtain their re-scaled versions $\{\mathbf{OI^{0.5}},\mathbf{BI^{0.5}}\}$ using auxiliary scaling factors. Subsequently, we input the multiscale image set $\{\mathbf{OI^{k}},\mathbf{BI^{k}}\},k\in\{0.5,1\}$ into ConvNeXt network. Then the deep features $\{\mathbf{OF^{k}},\mathbf{BF^{k}}\},k\in\{0.5,1\}$ are acquired through the split-transform-merge encoder strategy.
II-B Multiscale Feature Integration
The encoder features encompass a wealth of multi-scale information, where the details at various scales complement each other. Therefore, as presented in Fig. 3, we design the multiscale feature integration (MFI) module to fuse the interactive clues. Following [27], we firstly adopt interactively placed CBR Blocks, consisting of 3 × 3 convolutional layer, a batch normalization (BN) layer, and a ReLU activation function, to fuse the interactive clue, namely:
where $C(\cdot)$ and $f(\cdot)$ mean the concatenation operation along the channel dimension and the CBR block. $(\cdot)_{\uparrow}$ represents ×2 upsampling operation using bilinear interpolation.
Subsequently, we employ self-attention to further combine the features, opting for the widely-used Convolutional Block Attention Module (CBAM)[28] as our strategy. Finally, the features processed by the attention mechanism are fused with the input features via residual connections to supplement the information, producing the integrated feature vectors $\{\mathbf{OM^{0.5}},\mathbf{OM^{1}}\}$ through mean-pooling layers.
@@ II-C Multiscale Refinement Classifier
In the realm of image retrieval tasks, it has been demonstrated that the incorporation of a classifier significantly enhances performance[29, 30]. In this context, we have meticulously designed the multiscale refinement classifier (MRC), which is specifically tailored to refine the integrated features. Specifically, we use Linear Blocks, consisting of a linear layer, a batch normalization (BN) layer and a dropout layer, to obtain the refined feature representation$\{\mathbf{OR^{0.5}},\mathbf{OR^{1}}\}$. And then we obtain the fused class result $\mathbf{OC}$ by the classifier layers in OBI Branch. In addition, we apply the Kaiming initialization method[31] for the classifier component.
TABLE I: Quantitative comparison of retrieval success rate on the OBI-BI dataset. The best numbers are indicated in bold
Method Recall@1 Recall@5 Recall@10 Recall@top1 AP
@@
Ours-R 66.65 76.25 79.11 90.48 53.23
Ours-T 69.32 78.49 81.79 91.53 55.89
Ours-B 70.53 79.08 82.10 91.68 58.20
TABLE II: Quantitative result for the proposed CFIRN for different periods.
Method Recall@1 Recall@5 Recall@10 Recall@top1 AP
OBI→BI 69.32 78.49 81.79 91.53 55.89
@@ II-D Loss Function
Each ancient character is classified into a category of modern Chinese characters based on its developmental relationship, so the cross-font image retrieval problem can be transformed into an image classification problem. Specifically, we apply the widely used cross-entropy loss to optimize the parameters of the network, which can be represented:
@@
Through siamese network structure, the input images from the two branches are represented as multi-scale feature vectors. To reduce the distance between feature vectors of the same category, the conventional Euclidean distance-based triplet loss[32] is used, which can be formulated as:
@@
where $M$ represents the margin and set to 0.3 in all of experiments. $\|\cdot\|_{2}$ means the $L_{2}$-distance and $V_{i}$ is the feature vector of $i^{th}$ anchor image. $V_{p}$ and $V_{n}$ are the feature vectors of an image from the same and a different category as anchor respectively.
In summary, CFIRN can be optimized by minimizing the following losses:
@@
where $L$ means the total loss and the weight coefficient ${\alpha}$ is set to 5. $KL$ represents we also apply the KL divergence loss to enhance the training performance.
@@ III-A Experiment Setup
Datasets and Evaluation Criteria. We collect three ancient script image retrieval datasets for the purpose of OBI character decipherment, sourced from Anyang Normal University and the web. The gallery font include Bronze Inscription (BI), Bamboo Slip Inscription (BSI) and Clerical Script (CS). In the OBI-BI dataset, we conducted experiments with 1358 common character categories with 115655 images, with 30% of the categories allocated to testing, to simulate real-world challenges. And the same dataset partitioning strategy is applied to the OBI-BSI dataset and the OBI-CS dataset. Following studies in image retrieval, we employ several widely used metrics including Recall@K (R@K) and average precision (AP) to evaluate the performance of all models.
Implementation Details. Our proposed CFIRN is implemented with PyTorch [33], trained end-to-end for 340 epochs with a batch size of 8 on an NVIDIA 2080Ti GPU, and the ConvNeXt pretrained on ImageNet is employed as our vision backbone. The optimizer chosen is SGD with a momentum of 0.9 and a weight decay of 0.0005. During training, we resize the input image to 256 × 256 and apply image augmentation.
TABLE III: Ablation study of the components on the OBI-BI dataset. The best numbers are indicated in bold
MRC MFI Recall@1 Recall@5 Recall@10 Recall@top1 AP
@@
III-B Quantitative Evaluation
As an image retrieval problem across different domains, we compare our method with the state-of-the-art models, including S4G[25], MCCG[24] and RKNet[29], in similar cross-view image retrieval tasks. Table I highlights the significant advantages of our CFIRN in cross-font image retrieval, showcasing its superior performance. Specifically, ”Ours-R” and ”Ours-T/B” denote the backbone architectures of ResNet101 [34] and ConvNeXt-Tiny/Base, respectively.
In Table II, a series of experiments validate the model’s ability to effectively match characters from different historical periods. The OBI-BI scenario demonstrates notably better performance compared to others, closely reflecting real-world decipherment tasks. In essence, the closer the time periods, the greater the similarity in character shapes, leading to higher retrieval success.
III-C Qualitative Evaluation
This process vividly reveals the historical trajectory of Chinese character evolution and the intimate connection between character forms and their meanings. The comparative analysis results further confirm the accuracy of our experiments: characters closer in time to the OBI period are more similar in form to OBI, thus easier to match. This not only showcases the efficacy of our model but also provides a compelling basis for inferring the evolution from Oracle Bone Inscriptions to contemporary Chinese characters.
TABLE IV: Ablation study of different resolution strategy on the OBI-BI dataset. The best numbers are indicated in bold
Method Recall@1 Recall@5 Recall@10 Recall@top1 AP

### v2

@@ Abstract
Oracle Bone Inscription (OBI) is the earliest mature writing system in China, which represents a crucial stage in the development of hieroglyphs. Nevertheless, the substantial quantity of undeciphered OBI characters remains a significant challenge for scholars, while conventional methods of ancient script research are both time-consuming and labor-intensive. In this paper, we propose a cross-font image retrieval network (CFIRN) to decipher OBI characters by establishing associations between OBI characters and other script forms, simulating the interpretive behavior of paleography scholars. Concretely, our network employs a siamese framework to extract deep features from character images of various fonts, fully exploring structure clues with different resolutions by multiscale feature integration (MFI) module and multiscale refinement classifier (MRC). Extensive experiments on three challenging cross-font image retrieval datasets demonstrate that, given undeciphered OBI characters, our CFIRN can effectively achieve accurate matches with characters from other gallery fonts, thereby facilitating the deciphering.
@@ I Introduction
Oracle Bone Inscription (OBI) derives its name from being carved on tortoise shells and animal bones, serving as significant records for divination and military warfare during China’s Shang Dynasty. Therefore, deciphering OBI characters enables scholars to access valuable information recorded thousands of years ago, offering crucial insights into the social structure and cultural practices of early China. Specifically, as illustrated in Fig. 1(b), the process of deciphering involves identifying and correlating OBI characters with their corresponding modern Chinese counterparts. To decode undeciphered OBI characters, paleographers require not only extensive linguistic expertise and practical experience but also a deep understanding of the historical development and diverse variations of these characters [1]. Due to the labor-intensive and time-consuming nature of the interpretation process, only approximately 2,200 of the nearly 4,500 OBI characters discovered have been deciphered to date [2] as shown in Fig. 1(a).
@@
With significant advancements in artificial intelligence technology in recent years, an increasing number of researchers have applied advanced AI methods to address various practical challenges in the OBI domain, such as oracle bone character fragments rejoining [3, 4], OBI character detection [5, 6, 7, 8, 9], and OBI character recognition [10, 11, 12, 13, 14, 15, 16, 17, 18], to name a few.
Figure 1: (a) Over half of OBI characters remain undeciphered. (b) The goal of deciphering is to match OBI characters to corresponding modern Chinese characters. (c) Utilizing historical font intermediaries to overcome glyph gaps and enable effective decipherment of undeciphered OBI characters.
Figure 2: (a) Taking the OBI Branch as an example, high-quality feature extraction and class prediction are performed for the input image $\mathbf{OI^{1}}$. (b) The siamese network architecture of the proposed CFIRN. (c) The loss functions utilized in CFIRN include Cross-Entropy for improving classification accuracy, KL Divergence for aligning feature distributions, and Triplet Loss for optimizing the feature space.
@@
In the field of OBI character decipherment, significant progress have been made through various deep learning-driven approaches. Wang et al.[19] analyzed the evolution process using few-shot learning, while others have explored classification approaches to associate undeciphered OBI with ancient scripts such as the large seal script [20]. The use of diffusion models has also been emerged as a promising avenue for evolving OBI characters into modern forms [21, 22]. Furthermore, innovative methods involving the disassembly and reconstruction of OBI characters into modern substitutes have been proposed [23]. Gao et al.[24] proposed a retrieval-based method to establish a mapping relationship between undeciphered OBI characters and deciphered ones. Collectively, these studies underscore the potential of deep learning in advancing the automated decipherment.
@@
However, the significant evolution of character forms throughout an extensive historical period presents an impediment to establishing direct and immediate correlations. In the concrete practice of paleography, researchers often employ comparative analysis of OBI characters with scripts from various historical periods, particularly those from closely related epochs, to choose more plausible interpretations.
@@
Drawing insights from the methodology of paleographic research, we designed a workflow that utilizes gallery fonts from other historical periods as intermediaries to decipher the undeciphered OBI characters, as shown in Fig. 1(c). Concretely, we propose a cross-font image retrieval network (CFIRN) that establishes associations between OBI characters and other script forms, with the process depicted in Fig. 2. To address the dual challenge in ancient script images, where large-scale inputs emphasize intricate details and small-scale inputs capture broader structural patterns for complementary analysis, we seamlessly combine visual encoder features through the application of multiscale feature integration (MFI) module, further enhancing them with multiscale refinement classifier (MRC) to achieve the generation of highly refined features. Consequently, three richly annotated ancient scripts, including Bronze Inscription (BI), Bamboo Slip Inscription (BSI) and Clerical Script (CS), are selected as our gallery set. These scripts are essential due to their historical proximity to OBI and their role in capturing the structural transitions that lead to modern Chinese characters. Additionally, we conducted exploratory experiments on real-world undeciphered OBI data, where qualitative results further demonstrated the effectiveness of our model.
@@
In summary, our main contributions can be summarized as follows: 1) To the best of our knowledge, this is the first study to explore the capability of a unified cross-font image retrieval network for the task of OBI character decipherment. 2) Our CFIRN framework introduces a novel multiscale feature integration (MFI) module and a multiscale refinement classifier (MRC), to enhance multiscale feature extraction and fusion, enabling robust representation of ancient script characters. 3) Compared to state-of-the-art image retrieval models, our CFIRN demonstrates superior performance on three ancient script image retrieval datasets.
@@ II METHOD
The architecture of the proposed Cross-Font Image Retrieval Network (CFIRN), as shown in Fig. 2, leverages a Siamese network structure to integrate multiscale deep features extracted from two fonts, which enhances the quality of feature representation, simulating the interpretive behavior of comparing ancient inscriptions with known script forms. Both the OBI branch and the Gallery Font branch share the same framework, and for clarity, we describe the OBI branch as an example in Sections II-B and II-C.
II-A Feature Extraction Process
Drawing inspiration from related research [25, 26], we adopted ConvNeXt [27] as the backbone of the encoder network. The encoder weights are shared between the OBI branch and the Gallery Font branch, enabling better generalization in feature extraction under low-data conditions. Given an input image pair $\{\mathbf{OI^{1}},\mathbf{BI^{1}}\}$ (comprising an OBI character image and a corresponding gallery font image), we generate their scaled versions $\{\mathbf{OI^{0.5}},\mathbf{BI^{0.5}}\}$ using predefined scaling factors, where $\mathbf{OI^{k}}$ and $\mathbf{BI^{k}}$ represent the input images at scale $k$. These multiscale images $\{\mathbf{OI^{k}},\mathbf{BI^{k}}\},k\in\{0.5,1\}$ are then fed into the visual encoder. Using the split-transform-merge strategy, we obtain the deep features $\{\mathbf{OF^{k}},\mathbf{BF^{k}}\},k\in\{0.5,1\}$.
II-B Multiscale Feature Integration
The encoder features encompass a wealth of multi-scale information, where the intricate details at various scales intricately complement each other, capturing both fine-grained and coarse structural cues characteristic of ancient characters. To effectively combine these interactive clues, we propose the Multiscale Feature Integration (MFI) module, as illustrated in Fig. 3. Following the strategy in [28], we employ interactively placed CBR Blocks, each consisting of a 3×3 convolutional layer, a Batch Normalization (BN) layer, and a ReLU activation function, to fuse the multiscale features. Specifically, the fusion process is defined as follows:
where $C(\cdot)$ denotes channel-wise concatenation, $f(\cdot)$ represents the CBR block, and $(\cdot)_{\uparrow}$ indicates a ×2 upsampling operation using bilinear interpolation.
To further refine and combine the fused features, we incorporate the widely-adopted Convolutional Block Attention Module (CBAM), which applies both channel attention and spatial attention mechanisms. This enhances the feature representations by emphasizing important information while suppressing irrelevant details. Finally, the attention-enhanced features are integrated with the original input features via residual connections, ensuring information preservation. The integrated feature vectors $\{\mathbf{OM^{k}},\mathbf{BM^{k}}\},k\in\{0.5,1\}$ are then obtained using mean-pooling layers to produce compact and robust multiscale representations.
@@ II-C Multiscale Refinement Classifier
In the realm of image retrieval tasks, previous studies [29, 30] have demonstrated that incorporating a classifier can effectively improve retrieval performance. Here, we propose the Multiscale Refinement Classifier (MRC), which is specifically tailored to refine the integrated features. The MRC employs Linear Blocks, where each block consists of a linear layer, a Batch Normalization (BN) layer, and a dropout layer. These blocks transform the integrated features into the refined representations $\mathbf{OR^{k}},k\in\{0.5,1\}$, where each refined representation has a channel dimension of 512. The refined features are then passed through the classifier layers in the OBI Branch to produce the fused class prediction $\mathbf{OC}$. In the testing phase, the classifier is excluded, and the refined representations $\mathbf{OR^{k}},k\in\{0.5,1\}$ are concatenated to serve as the final feature representation for OBI characters. Additionally, we initialize the classifier components using the Kaiming initialization method to ensure stable and efficient training.
TABLE I: Overview of Collected Font Image Datasets from Four Historical Periods, Categorized by Modern Chinese Character Classes
Font #images #classes
Oracle Bone Inscription (OBI) 74994 1981
Bronze Inscription (BI) 66279 3354
Bamboo Slip Inscription (BSI) 18493 775
Clerical Script (CS) 29596 5250
TABLE II: Quantitative comparison of retrieval success rate on the OBI-BI dataset. The best numbers are indicated in bold
Method Recall@1 Recall@5 Recall@10 Recall@1% AP
@@
Ours-R50 63.09 73.75 77.15 89.39 49.33
Ours-R101 66.65 76.25 79.11 90.48 53.23
Ours-C-T 69.32 78.49 81.79 91.53 55.89
Ours-C-S 70.52 78.79 81.36 91.10 58.27
Ours-C-B 70.53 79.08 82.10 91.68 58.20
TABLE III: Quantitative result for the proposed CFIRN for different periods.
Method Recall@1 Recall@5 Recall@10 Recall@1% AP
OBI→BI 70.53 79.08 82.10 91.68 58.20
@@ II-D Loss Function
Each ancient character is assigned to a corresponding category of modern Chinese characters based on their developmental relationships. As a result, the cross-font image retrieval task is reformulated as an image classification problem, as illustrated in Fig. 2(c). To optimize the network parameters, we adopt the widely-used cross-entropy loss $CEL_{i}$, which measures the discrepancy between the predicted class distribution and the ground truth. The formulation is as follows:
@@
Through siamese network structure, the input images from the two branches are represented as multi-scale feature vectors. To align the feature vectors of images from the same category while separating those from different categories, we utilize the triplet loss [31], which is defined as:
@@
where $M$ is the margin value, set to 0.3 in all experiments, and $\|\cdot\|_{2}$ denotes the $L_{2}$-distance. Here, $V_{i}$ is the feature vector of the anchor image, $V_{p}$ is the feature vector of a positive sample (same category), and $V_{n}$ is the feature vector of a negative sample (different category).
In addition, to ensure consistency between the predictions of the OBI Branch and the Gallery Font Branch, we employ the Kullback-Leibler (KL) divergence loss. The KL loss aligns the probability distributions of the two branches, which helps to improve generalization and stabilize the training process for cross-font retrieval tasks.
In summary, the overall loss function for CFIRN combines three components: cross-entropy loss $CEL$, KL divergence loss $KL$, and triplet loss $TL$. The total loss $L$ is formulated as::
@@
where the weight coefficient ${\alpha}$ is set to 5. This combination ensures robust feature representation and improved retrieval performance.
@@ III-A Experiment Setup
Datasets and Evaluation Criteria. As shown in Table I, we collect three ancient script image retrieval datasets for the purpose of OBI character decipherment, sourced from Anyang Normal University and the web. The gallery font include Bronze Inscription (BI), Bamboo Slip Inscription (BSI) and Clerical Script (CS). In the OBI-BI dataset, we conducted experiments with 1358 common character categories with 115655 images, with 30% of the categories allocated to testing, to simulate real-world challenges. And the same dataset partitioning strategy is applied to the OBI-BSI dataset and the OBI-CS dataset. Following studies in image retrieval, we employ several widely used metrics including Recall@K, Recall@Q%, and average precision (AP) to evaluate the performance of all models.
Implementation Details. Our proposed CFIRN is implemented with PyTorch, trained end-to-end for 340 epochs with a batch size of 8 on an NVIDIA 2080Ti GPU, and the ConvNeXt pretrained on ImageNet is employed as our vision backbone. Learning rate initialization is 0.0015 for the backbone parameters and 0.005 for the other layers. The warm-up period lasts for 80 epochs, after which the learning rate is reduced by a factor of 0.1 at 180 and 240 epochs. The optimizer chosen is SGD with a momentum of 0.9 and a weight decay of 0.0005. During the training process, the input images are resized to 256 × 256, and several image augmentation methods are applied, such as random padding, random cropping, and random flipping.
TABLE IV: Ablation study of the components on the OBI-BI dataset. The best numbers are indicated in bold
MRC MFI Recall@1 Recall@5 Recall@10 Recall@1% AP
@@
TABLE V: Ablation study of different resolution strategy on the OBI-BI dataset. The best numbers are indicated in bold
Method Recall@1 Recall@5 Recall@10 Recall@1% AP
@@
III-B Quantitative Evaluation
As an image retrieval problem across different domains, we compare our method with the state-of-the-art models, including S4G[26], MCCG[25] and RKNet[29], in similar cross-view image retrieval tasks. Table II highlights the significant advantages of our CFIRN in cross-font image retrieval, showing its superior performance. Specifically, ”Ours-R50/R101” and ”Ours-C-T/S/B” denote the backbone architectures of ResNet-50/101 [32] and ConvNeXt(Tiny/Small/Base), respectively. ”Ours-B” achieves the best performance with a Recall@1 of 70.53% and an AP of 58.20%, surpassing the state-of-the-art RKNet by 9.55% in Recall@1 (from 64.39% to 70.53%) and 7.00% in AP (from 51.20% to 58.20%).
In Table III, a series of experiments validate the model’s ability to effectively match characters from different historical periods. The OBI→BI scenario achieves the best performance with a Recall@1 of 70.53% and an AP of 58.20%, significantly outperforming the other two scenarios. This reflects the closer historical proximity and greater structural similarity between Oracle Bone and Bronze Inscriptions, leading to higher retrieval success. In comparison, the OBI→BSI task achieves moderate performance (Recall@1: 42.68%, AP: 32.52%), while the OBI→CS task shows the lowest scores (Recall@1: 34.71%, AP: 28.85%),demonstrating the increased difficulty as the time gap widens. These results align well with real-world paleographic decipherment, where characters from closer historical periods retain more structural similarities, making them easier to associate. The findings further validate the robustness of CFIRN in handling cross-font image retrieval tasks across varying levels of character evolution.
III-C Qualitative Evaluation
The evolution of Chinese characters exhibits diverse morphological variations over time. For instance, certain smaller components may enlarge (Fig. 4, line 1), shift their spatial positions (Fig. 4, line 2), or undergo changes in stroke structure (Fig. 4, line 3). Additionally, the overall character shape may experience left-right flipping (Fig. 4, line 4), further increasing the complexity of character transformations.
As demonstrated by a series of challenging samples, our model successfully retrieves the most similar characters from the gallery datasets, achieving precise matches (highlighted in yellow) across different fonts. These correctly matched results not only reflect the robustness of CFIRN in handling character variations over time but also highlight its ability to bridge the morphological gap between OBI characters and their modern Chinese counterparts.